40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 7 20:32:21 2009, rana, Update, Adaptive Filtering, extra delay and noise in PEM -> ASS/OAF system a.gif
    Reply  Mon Oct 19 11:31:55 2009, Jenne, Update, Adaptive Filtering, extra delay and noise in PEM -> ASS/OAF system PEM-timing_19Oct2009.png
       Reply  Mon Oct 19 19:37:39 2009, Sanjit, Jenne, Update, Adaptive Filtering, extra delay and noise in PEM -> ASS/OAF system PEM_timingDealy_19OCT09_MCL2PEM2.png
          Reply  Tue Oct 20 11:38:10 2009, rana, rolf, Update, Adaptive Filtering, extra delay and noise in PEM -> ASS/OAF system 
Message ID: 2116     Entry time: Mon Oct 19 11:31:55 2009     In reply to: 2066     Reply to this: 2121
Author: Jenne 
Type: Update 
Category: Adaptive Filtering 
Subject: extra delay and noise in PEM -> ASS/OAF system 

Quote:

[Rana, Jenne]

There is some craziness going on with the delay in the PEM path for the OAF.  We plot the difference between the C1:PEM-SEIS_GUR1_X and C1:ASS-TOP_PEM_10.  These are physically the same channel, plugged into the PEM ADCU, and then the signal is used as a regular PEM channel, and is also sent to the ASS computer and used there for the OAF system.  As you can see in the blue trace on the bottom plot, there is a huge amount of delay, and it's very noisy.  We also plot the _GUR2_X / ASS-TOP_PEM_2 pair (red), and it has a similar amount of delay, but it is not nearly as fuzzy and noisy.  For comparison, we plot the SUS-MC2_MCL (which is identical to IOO-MC_L) and ASS-TOP_ERR_MCL pair (green), and they don't have any big overall delay problems, so it's not totally a problem with the signals getting to the ASS computer.

This problem was present during/after all of the following attempts to fix it:

* The sample rate on the ASS computer is 2048.  The PEM channels were being acquired the ADCU at 512.  We changed the ADCU sampling rate to 2048 to match.

* We soft rebooted the ASS computer, in case it was a timing problem.

* Doing a "sudo shutdown -r now" while logged in as controls.

We might also try resetting/power cycling c0dcu in the morning.  Alex has been emailed to help us try to figure this out.

 

In other news, the time delay that we measure from the plot gives us 180degrees in ~210Hz.  This corresponds to a little more than 2msec of delay, with the C1:ASS version lagging behind the C1:PEM version.  (2 samples at 840Hz) Converting to the 2048 sampling rate, we have a delay of 4.8, so 5 front-end cycles.  Since Rana measured this morning that the delay indicated by the transfer function is 10 cycles, and this delay shows that the ASS lags the actual seismometer signal by 5 cycles, we should subtract this 5 from the 10 from the transfer function, giving us a final sample-and-hold delay of 5.  Coincidentally(?), 5 is the delay that was found in the C1:ASS-TOP screen, after it's one year of dormancy.  The point of the delay feature in the code is to help match the delay in the two signal paths: the PEM path and the output path of the filter.  Since the output has a lag of 10, and the PEM path has a lag of 5, to make them match, we artificially put in a delay of 5.

 Alex came in a week ago Friday to help figure this timing problem out, and some progress was made, although there's more to be done. 

Here are the (meager) notes that I took while he was working:

we can rename the tpchn_C1_new back to tpchn_C1, but the _new one works right now, so why change it?

need to find dcuDma.c source code...this is (?) what sends the PEM channels over to ASS.  Found:  source code is dcu.c, th
en the binary is dcuDma.o  Trying to recompile/remake dcuDma to make everything (maybe) good again.

Possibility: maybe having so many channels written to the RFM takes too long? shouldn't be  a problem, but maybe it is.  I
n the startup.cmd (or similar?) change the number of ISC modules to 1, instead of 2, since we only have one physical board
 to plug BNCs into, even though we have 2 isc boards.  c0dcu1 rebooted fine with the one isc board.  now can't get ass tes
tpoints to try the DTT timing measurement again.  rebooting fb40m to see if that helps.  fb40m is back, but we still don't
 have ASS testpoints.  Alex had to leave suddenly, so maybe more later.

Also, next possibility is that c0dcu and c1ass are not synched together properly....we should look at the timing of the AS
S machine.

 

After these adventures, the noisy trace in the timing delay (in the plot in elog 2066) has become quiet, as shown below (The blue trace, which was noisy in 2066 is now hiding behind the red trace).  However, the overall timing delay problem still exists, and we don't quite understand it.  Alex and I are meeting tomorrow morning at the 40m to try and suss this out.  Our first plan of attack is to look at the ASS code, to see if it puts any weird delays in.

Attachment 1: PEM-timing_19Oct2009.png  75 kB  | Hide | Hide all
PEM-timing_19Oct2009.png
ELOG V3.1.3-